A manufacturing company forecast that it is likely to sell 8, 00,000 units for the year 2021. The processing cost of an order is Rs.200 and the carrying cost per unit of inventory is Rs.16. The lead time of an order is 5 days.(a) What would be the economic order quantity (EOQ) and re-order point assuming 360 days in a year? (5 Marks)(b) The company implements business process reengineering which results in to reduction of 25% in cost of an order, 15% in carrying cost per unit of inventory and 20% in lead time of an order. What would be the new EOQ and re-order point. 

1
Expert's answer
2021-08-29T16:52:44-0400

Solution:

a.). The Economic Order Quantity (EOQ) = "\sqrt{\frac{2DS}{H} }"

Where: D = Annual quantity demanded = 800,000

            S = Ordering cost = 200

            H = Holding cost per unit = 16


The Economic Order Quantity (EOQ) = "\sqrt{\frac{2\times 800,000 \times 200}{16} }" = "\sqrt{20,000,000 } = 4,472 \;units" 

The Economic Order Quantity (EOQ) = "4,472 \;units"


Re-order point = Average daily unit sales "\times" lead time


Where: Average daily unit sales = "\frac{800,000}{360} = 2,222"

            Lead time = 5 days

Re-order point = 2222 "\times" 5 = 11,110 units


b.). New cost of an order = 200 – (200 "\times" 25"\%" ) = 200 – 50 = 150

New carrying cost = 16 – (16 "\times 15\%") = 16 – 2.4 = 13.6


The Economic Order Quantity (EOQ) = "\sqrt{\frac{2DS}{H} }"

Where: D = Annual quantity demanded = 800,000

            S = Ordering cost = 150

            H = Holding cost per unit = 13.6


The Economic Order Quantity (EOQ) = "\sqrt{\frac{2\times 800,000 \times 150}{13.6} } = \sqrt{17,647,059 } = 4,200 \; units"

The new Economic Order Quantity (EOQ) = 4,200 units

 

Re-order point = Average daily unit sales "\times" lead time


Where: Average daily unit sales = "\frac{800,000}{360} = 2,222"

            Lead time = 5 – (5 "\times 20\%") = 5 – 1 = 4 days


Re-order point = 2222 "\times" 4 = 8,888 units


New Re-order point = 8,888 units